June 25, 2010 June 25, 2009 FX per $ FX per £ 1.721 1.559 9.045 1.215 11.643 69.476 FX per 1.417 1.283 7.443 1.000 9.583 57.179 110.308 15.631 9.577 0.822 1.232 FX per $ 1.225 1.084 5.238 0.703 7.750 Country (currency) Australia (dollar) Canada (dollar) Denmark (krone) Euro Hong Kong (dollar) India (rupee) Japan (yen) Mexico (peso) Sweden (krona) United Kingdom (pound) United States (dollar) Source: U.S. Federal Reserve Board of Governors, H.10 release: Foreign Exchange Rates. 1.152 1.037 6.036 0.811 7.779 46.36 89.35 48.16 94.86 13.22 134.048 12.697 7.74 0.667 1.000 18.993 11.632 1.000 1.496 7.460 0.609 1.000

2. Using the information in the table for June 25, 2010, compute the Danish krone-Canadian dollar exchange rate Etrone/C (a) 0.172 (b) 1.037 (c) 5.821 (d) 6.036 3. If EvsD/CAD is 0.81 and EusD/GBP is 1.56 what must be EGBP/CAD? (a) 0.52 (b) 1.26 (c) 1.93 (d) 2.37
Answer #1

2. In June 25, 2010

EC$/USD = 1.037

EKrone/ USD = 6.036

Therefore, Ekrone/ C$ = (6.036/1.037) = 5.821

Hence, option(C) is correct.

3. EUSD/CAD = 0.81

EUSD/GBP =1.56

Therefore, EGBP/CAD = (0.81/1.56) = 0.52

Hence, option(A) is correct.